OpenStudy (anonymous):

1. Solve 6C3 A. 120 B. 720 C. 20* D. 240 2. Which is not a combination? A. Choosing 3 topping for your pizza B. Lining 3 students up in a row* C. Choosing 2 deserts from a tray of 10 D. Choosing 5 students to represent a class of 30 3. How many ways can you choose 3 dishes from a menu of 8? A. 24 B. 336 C. 144 D. 56* 4. Is choosing 5 players to represent the team at a fund raiser a permutation or combination? A. Permutation B. Combination* 5. Solve 6P4 A. 240 B. 720 C. 20 D. 360*

OpenStudy (anonymous):

6. How many different ways could you arrange 4 books on a shelf? A. 4 B. 24* C. 14 D. 8 7. Is choosing two toppings from a list of 8 for your pizza a permutation or combination? A. Permutation B. Combination* 8. How many different ways can you choose 3 frozen yogurt toppings from 20? A. 6,840 B. 1,140* C. 60 D. 2,260

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (amistre64):

none of this seems appropriate ...

OpenStudy (amistre64):

what context am i missing to justify this kind of posting?

OpenStudy (anonymous):

This is my answers, I just want someone to check them.

OpenStudy (anonymous):

B D B D B B B 8C3 8! / [3!(8 - 3)!] 8! / (3! * 5!) (8 * 7 * 6 * 5 * 4 * 3 * 2) / (3 * 2 * 5 * 4 * 3 * 2) (8 * 7 * 6) / (3 * 2) 8 * 7 56 6p4 6! / (6 - 4)! 6! / 2! (6 * 5 * 4 * 3 * 2) / 2 360 4! = 4 * 3 * 2 = 24 20C3 = 20! / [3!(20 - 3)!] 20! / (3! * 17!) (20 * 19 * 18) / (3 * 2) 1140

OpenStudy (amistre64):

answer checking ... thats better than what it looked like to start with

OpenStudy (anonymous):

This is how you set the problems up and solve them.

OpenStudy (anonymous):

Thank you! :D This helps

OpenStudy (anonymous):

Got them ALL right again. :D